A wheel is rotating about an axis that is in the z -direction. The angular velocity ω z is -6.00 rad/s at t = 0, increases linearly with time, and is +4.00 rad/s at t = 19.0 s . We have taken counterclockwise rotation to be positive.

Part B

During what time interval is the speed of the wheel increasing? Express your answer with the appropriate units.

Part C

During what time interval is the speed of the wheel decreasing?

Part D

What is the angular displacement of the wheel at t = 19.0 s ?
Answer #1

A)

Angular acceleration, a = (4 + 6)/19 = 0.53 rad/sec^2

B)

6 = 0.53 t

Time, t = 11.4 sec

C)

4 = 0.53 t

Time, t = 7.55 sec

D)

Theta = (-6 x 19) + (0.5 x 0.53 x 19^2) = - 18.34 rad

Comment in case any doubt please rate my answer...